# Generative AI in Sociological Research:

### v1

@@ Abstract
Forthcoming, Sociological Science
Generative artificial intelligence (GenAI) has garnered considerable attention for its potential utility in research and scholarship, even among those who typically do not rely on computational tools. Early commentators, however, have also articulated concerns about how GenAI usage comes with enormous environmental costs, serious social risks, and a tendency to produce low-quality content. In the midst of both excitement and skepticism, it is crucial to take stock of how GenAI is actually being used. Our study focuses on sociological research as our site, and here we present findings from a survey of 433 authors of articles published in 50 sociology journals in the last five years. The survey provides an overview of the state of the discipline with regard to the use of GenAI by providing answers to fundamental questions: how (much) do scholars use the technology for their research; what are their reasons for using it; and how concerned, trustful, and optimistic are they about the technology? Of the approximately one third ofrespondents who self-report using GenAI at least weekly, the primary uses are for writing assistance and comparatively less so in planning, data collection, or data analysis. In both use and attitudes, there are surprisingly few differences between self-identified computational and non-computational researchers. Generally, respondents are very concerned about the social and environmental consequences of GenAI. Trust in GenAI outputs is low, regardless of expertise or frequency of use. While optimism that GenAI will improve is high, scholars are divided on whether GenAI will have a positive impact on the field.
keywords:
generative AI; sociology of science; large language models; sociological research practices; computational sociology; survey
Reproducibility Package: A replication repository for this paper can be found at: https://github.com/Marshall-Soc/genai_sociology. The data for this paper is hosted on the Harvard Dataverse [data_repo], and can be accessed through: https://doi.org/10.7910/DVN/ICXIRP
@@ 1 Introduction
Thus far, the use of GenAI as a tool for sociological research has mainly been explored from the perspective of research products. Research articles offer “proof of concepts” by identifying important use cases and limitations (see, e.g., [davidskar24, Boelaert2025-la]). Meanwhile, less attention has been given to how and why scholars in our discipline decide to use or avoid GenAI [c.f. Watermeyer2024-kp, Watermeyer2024-pt]. Additionally, while there have been a series of recent publications that showcase high-level applications of GenAI—say, classifying images [Law2024-jn] or extracting information from text [Stuhler2025-rp]—there has been comparatively little open discussion about how sociologists may use GenAI for more every-day aspects of research – be it for writing, generating ideas, debugging code, or otherwise.
In this paper, we present a survey for which we asked sociologists and their collaborators about how they perceive and use (or do not use) GenAI in their research. Our sampling frame includes the authors of all articles published in 50 sociology journals in the last five years. We contacted a random sample of these authors, as well as all authors who published ’computational’ articles. This allows us to paint the landscape of GenAI usage and attitudes among the population of researchers contributing to recent sociological scholarship. Going beyond speculation about what technology “could” do for sociologists, we offer the first systematic evidence on how GenAI is currently being used and seen by sociologists. Thereby, we provide an empirical baseline for a more grounded discussion on GenAI’s role in sociological research.
@@ 2.1 What Is “Generative” AI?
Language surrounding “artificial intelligence” is often imprecise, with a wide range of technologies grouped under this umbrella [Bender2025-wa, 1-20]. Therefore, we begin by briefly explaining how we define GenAI in this study (also see our primer on GenAI in email invitation sent to survey respondents in A).
The distinguishing characteristic of GenAI from other computational methods is in the name: as opposed to “discriminative” models that find optimal boundaries in data, primarily for classification tasks, generative models are designed to generate text, images, audio, and video. We focus on text generation models, as that is also the focus of the current literature in sociology and represents the most common form of GenAI interaction [Zhang2025-st]. When employing such models, users typically write a prompt containing the text that a trained model will take as its point of departure, such as instructions, questions, or an arbitrary query. This prompt is encoded as tokens [^1] Tokens are the basic inputs and outputs in natural language processing. In LLMs, tokens are words, combinations of words (including complete sentences), parts of words (e.g., word stems), and/or common non-word characters like spaces or dashes. See [mielke2021between] for additional detail., then the underlying model generates a probable continuation of the prompt by identifying the most likely next token, defined in probabilistic terms. The model then incorporates that new token into another round of generation and continues until a stopping rule, such as a maximum number of tokens to generate, is satisfied. In other words, the prompt provides a starting point from which the model starts generating ([for a more thorough introduction and discussion of practical aspects, we recommend Chae2023-yb]).
@@ 2.3 Generative AI in Sociological Research
[…]
Sociologists have also pointed out how GenAI complicates sociological analysis. This includes homogenization of outputs [Zhang2025-st, Alvero2024-iq]; injecting new forms of machine bias [Boelaert2025-la, Stuhler2025-rp, Rister-Portinari-Maranca2025-dz]; and LLM’s parotting of government-approved ideological positions [Waight2024-tj] or potentially the positions favored by the organizations producing the models [Martin2023-mk]. In some cases, GenAI may perform only on par with, or even worse than, traditional methods that are often more transparent and less computationally costly [Mu2023-rg, Nelson2025-zb, Stuhler2025-rp, Ashwin2025-af]. Furthermore, if sources of empirical data, such as open-ended survey responses [Zhang2025-st] or social media posts [sourati2025shrinking] are increasingly produced using GenAI, future research will be confronted with difficult questions about sociological explanation and inference.
@@ 3.1 Sample
We used a bibliometric multistage sampling design to construct a representative sample of both computational and non-computational sociologists and their collaborators. The sample includes 219 non-computational authors and an over-sample of 214 computational authors (defined as those authoring an article with computational terms in the title, abstract, and keywords) for a total of 433 respondents. We used rake weights [debell2009computing] to adjust for this oversampling as well as selection bias into our sample—namely, gender and location. Our sampling method is detailed in B. Figure 1 is a summary schematic of the sampling strategy.
@@ 3.2 Measures
Our survey focuses on who uses GenAI, its use cases, reasons for and concerns over using GenAI, and sources of optimism and trust in the technology. See the replication repository2 for information about accessing the survey questionnaire and data, which includes variables not explored in the main text.
@@ 3.3 Analytic Plan
One particular interest of our work is to describe differences between scholars who self-report using computational methods and those who do not. To do this, we use an array of univariate, bivariate, and multivariate descriptive data visualizations. We also use regression models to analyze the extent trust and optimism are associated with levels of GenAI use and familiarity. All descriptive analyses and regressions are rake-weighted with weight-adjusted standard errors and confidence intervals [lumley2017fitting]. See C for our findings as percentages over individual respondent categories (e.g., $X\%$ of respondents stated they “somewhat agree” with $Y$). Recall that our sample includes sociologists and their collaborators. To simplify, we refer to this population collectively as scholars.
@@ 4.2 GenAI Use Cases
Of those who have used GenAI in their research at least once, how do they use it? The second panel of Figure 2 shows the distribution of GenAI use at various stages of research—planning, writing, data collection, and analysis—separated by non-computational and computational scholars. The percentages sum to 100% per research stage, but with the never-users included in the calculation and removed from the visualization. As the bar chart shows, computational scholars are slightly more likely to use GenAI in analysis tasks (21.9%; 16% $\leq\mu\leq$ 29.2%) than non-computational scholars (13.6%; 8.4% $\leq\mu\leq$ 21.4%). Outside this research stage, however, the differences are marginal. Importantly, while the never-users are not shown, they make up the largest group for each research stage—in other words, most scholars do not use GenAI in any of the research stages as defined by the survey. Nonetheless, among those who have used GenAI, writing is the most common research stage where GenAI comes into play for both non-computational and computational scholars. In the Appendix (Figure 5 in C) we break down each research stage into specific tasks, such as translating one’s own writing or explaining statistical outputs.  [^2] Generally, we find very few differences, the most significant of which is in the analysis stage. Specifically and unsurprisingly, computational scholars are using GenAI to help with coding – writing (20.8%) and debugging (21.8%). Non-computational scholars are also somewhat more likely to use GenAI for grammar and spelling assistance (32.6%) and translation (24.8%). That said, both computational and non-computational scholars most often use GenAI to help with writing tasks. Though analysis is tied with writing for computational scholars
@@ 4.4 Concerns about GenAI
Overall, our items indicate that there is a high level of concern about various aspects of GenAI; of the nine questions we asked, seven were rated as being between “Concerned” and “Very Concerned” on average. To capture possible nuance regarding respondent’s perceptions of risks, we also asked about their GenAI concerns in an open-ended question. For instance, one respondent wrote that they were “wholly against the use of GenAI tools. They have no place in social science practice.” Below, we complement the aggregate findings with further examples that we found to be representative of general themes in the open-ended responses.
@@ 4.5 Optimism and Trust
Finally, we measure the optimism and trust about the future of GenAI for research. Scholars are generally more optimistic that GenAI will continue to improve over the next 2 to 3 years (40.2% of all scholars “Somewhat agree” or “Strongly agree”)—though there is ambivalence about whether this will have a net positive effect on the field or whether the current advantages of GenAI outweigh the drawbacks for the field. Furthermore, scholars generally believed that GenAI outputs cannot be trusted (only 4.5% of all scholars “Somewhat agree” or “Strongly agree” that GenAI can be trusted, see also Figure 10 in C).
@@
I treat GenAI as an inexperienced, mediocre RA. I don’t trust it, but use it to do simple things I can easily check” and “I think we’re at the early stages here. I think it can sometimes be useful to bounce ideas off, but often it returns quite derivative feedback, and I have not found these tools very useful for searching for articles.
We used rake-weighted generalized linear regression models to further probe heterogeneity in trust and optimism. [^3] We also ran a series of rake-weighted ordered logistic regression models as a robustness check given the ordinal nature of the outcome variables. They are consistent with the simpler linear model results reported here. These alternative models are in Table 6 in E. Adjusted predictions for these models are visualized in Figure 11. Specifically, we examine whether scholars who are more familiar with GenAI are more likely to trust GenAI, believe that GenAI will improve in the next 2 to 3 years, or believe that GenAI will have a net positive impact on their field in the next 2 to 3 years, and if those effects might be moderated by their use level.
@@
The regression results are in Table 5 in D with unstandardized estimates. A model with only familiarity as a predictor of trust (not shown in 5) had a very small positive effect ($\hat{\beta}$ = 0.078 with $t$ = 3.660). However, use level is not a significant moderator of this effect (see Model 1 in Table 5). Use level is also not a significant moderator for the net positive outcome variable but is for the future improvement variable. The regression results are summarized with adjusted predictions in Figure 4, where gender identity is held at its modal value (cis-man).
@@ 5 Discussion
Second, we do not yet know to what extent sociologists might see GenAI as a threat to their sociological expertise. Sociological work may be challenged by GenAI as even the specialties most removed from computational methodology, namely sociological theory, are implicated. [^5] Incidentally, the most elite sociology departments in the US specialize in theory [elder2025status]. More generally, GenAI threatens knowledge work by broadening the availability of required expertise, weakening occupational closure mechanisms, and reducing bargaining power. Such tensions are further exacerbated by the increased dependence on science and technology in tandem with a decreasing trust in scientists [eyal2019crisis]. This routinization of knowledge production may eventually allow for the displacement of well-paid professionals for “workers paid a fraction of their wage, flood[ing] the market with inferior products, and impos[ing] unreasonable and punishing working conditions on those” using GenAI ([Bender2025-wa, 39], c.f., [Nelson2025-zb]). Social scientists might come to feel these tensions more acutely than they do now, given that GenAI can present convincing or even compelling arguments about human society and social behavior, even if these are inaccurate or entirely fabricated.
@@
[…]
Appendix A Example Recruitment Materials
A.1 Invitation Email
@@ A.2 Generative AI Explainer
Appendix B Sampling Methods
B.1 Design
A scholar’s expertise [contributory or interactional à la collins2019rethinking] in relation to GenAI is likely to shape both usage and attitudes. Specifically, we expected that computational scholars would bring a potentially unique perspective on GenAI as it has emerged out of the same computational techniques that many such scholars use (i.e., machine learning and NLP). However, comparatively, this group is a very small subset of the broader sociological community. We therefore make an effort to both identify these scholars from our larger sampling frame and oversample them. In addition to having a better understanding of how these technologies operate, we also expected that computational scholars will have a higher response rate due to intrinsic interest and will be more likely to use GenAI in some capacity. Furthermore, based on prior surveys of (generative and non-generative) AI usage in academia, we expected that men will also be more likely to respond (possibly, in part, because they are overrepresented among computational scholars). We therefore design the survey to allow us to adjust for some parameters of our population. Although we attempt to control for such bias, inferences to the entire field of sociology must be made with caution. See the schematic representation of our sampling strategy in the main text(Figure 1).

### v2

@@ Abstract
Generative artificial intelligence (GenAI) has garnered considerable attention for its potential utility in research and scholarship. A growing body of work in sociology and related fields demonstrates both the potential advantages and risks of GenAI, but these studies are largely proof-of-concept or specific audits of models and products. We know comparatively little about how sociologists actually use GenAI in their research practices and how they view its present and future role in the discipline. In this paper, we describe the current landscape of GenAI use in sociological research based on a survey of authors in 50 sociology journals. Our sample includes both computational sociologists and non-computational sociologists and their collaborators. We find that sociologists primarily use GenAI to assist with writing tasks: revising, summarizing, editing, and translating their own work. Respondents report that GenAI saves time and that they are curious about its capabilities, but they do not currently feel strong institutional or field-level pressure to adopt it. Overall, respondents are wary of GenAI’s social and environmental impacts and express low levels of trust in its outputs, but many believe that GenAI tools will improve over the next several years. We do not find large differences between computational and non-computational scholars in terms of GenAI use, attitudes, and concern; nor do we find strong patterns by familiarity or frequency of use. We discuss what these findings suggest about the future of GenAI in sociology and highlight challenges for developing shared norms around its use in research practice.
@@ 1 Introduction
Thus far, the use of GenAI as a tool for sociological research has mainly been explored from the perspective of research products. Research articles offer “proof of concepts” by identifying important use cases and limitations (see, e.g., [davidskar24, Boelaert2025-la]). Meanwhile, less attention has been given to how and why scholars in our discipline decide to use or avoid GenAI [Watermeyer2024-kp, Watermeyer2024-pt]. Additionally, while there have been a series of recent publications that showcase high-level applications of GenAI—say, classifying images [Law2024-jn] or extracting information from text [Stuhler2025-rp]—there has been comparatively little open discussion about how sociologists may use GenAI for more every-day aspects of research – be it for writing, generating ideas, debugging code, or otherwise.
In this paper, we present a survey for which we asked sociologists and their collaborators about how they perceive and use (or do not use) GenAI in their research. Our sampling frame includes the authors of all articles published in 50 sociology journals in the last five years. We contacted a random sample of these authors, as well as all authors who published “computational” articles. This allows us to paint the landscape of GenAI usage and attitudes among the population of researchers contributing to recent sociological scholarship. Going beyond speculation about what technology “could” do for sociologists, we offer the first systematic evidence on how GenAI is currently being used and seen by sociologists. Thereby, we provide an empirical baseline for a more grounded discussion on GenAI’s role in sociological research.
@@ 2.1 What Is “Generative” AI?
Language surrounding “artificial intelligence” is often imprecise, with a wide range of technologies grouped under this umbrella [Bender2025-wa]. Therefore, we begin by briefly explaining how we define GenAI in this study (also see our primer on GenAI in email invitation sent to survey respondents in Appendix B).
The distinguishing characteristic of GenAI from other computational methods is in the name: as opposed to “discriminative” models that find optimal boundaries in data, primarily for classification tasks, generative models are designed to generate text, images, audio, and video. We focus on text generation models, as that is also the focus of the current literature in sociology and represents the most common form of GenAI interaction [Zhang2025-st]. When employing such models, users typically write a prompt containing the text that a trained model will take as its point of departure, such as instructions, questions, or an arbitrary query. This prompt is encoded as tokens [^1] Tokens are the basic inputs and outputs in natural language processing. In LLMs, tokens are words, combinations of words (including complete sentences), parts of words (e.g., word stems), and/or common non-word characters like spaces or dashes. See [mielke2021between] for additional detail., then the underlying model generates a probable continuation of the prompt by identifying the most likely next token, defined in probabilistic terms. The model then incorporates that new token into another round of generation and continues until a stopping rule, such as a maximum number of tokens to generate, is satisfied. In other words, the prompt provides a starting point from which the model starts generating; for a more thorough introduction and discussion of practical aspects, see [Chae2023-yb].
@@ 2.3 Generative AI in Sociological Research
Sociologists have also begun to point out potential use cases of GenAI models for research tasks [Bail2024-pa, Davidson2024-mo]. Empirical studies have examined whether GenAI can increase research efficiency, for example, by generating survey questions [Gotz2024-lf]; imputing missing data [Kim2023-fn]; engaging in an exploratory dialogue with qualitative data [Ibrahim2024-ml, Hayes2025-bn]; and classifying, annotating, and extracting information from text or images [Gilardi2023-re, Law2024-jn, Stuhler2025-rp, Nelson2025-zb, Maranca2025-li, Schwitter2025-li, Lin2025-vk]. Another branch of scholarship has explored GenAI’s potential for simulating human behavior [Kozlowski2025-qe, Anthis2025-xh, Broska2024-od, Alvero2024-iq]. Much of this work has focused on simulating responses to survey items [Boelaert2025-la, Kim2023-fn, Broska2025SMR, Kozlowski2024-co, Zhang2025-st]; outside of sociology, see, e.g., [Argyle_2023], but GenAI can also be used to simulate human interaction and communication [Karell2024-qn, Horton2023-do, argyle2023leveraging].
Sociologists have also pointed out how GenAI complicates sociological analysis. This includes homogenization of outputs [Zhang2025-st, Alvero2024-iq]; injecting new forms of machine bias [Boelaert2025-la, Stuhler2025-rp, Rister-Portinari-Maranca2025-dz]; and LLMs parroting government-approved ideological positions [Waight2024-tj] or potentially the positions favored by the organizations producing the models [Martin2023-mk]. In some cases, GenAI may perform only on par with, or even worse than, traditional methods that are often more transparent and less computationally costly [Mu2023-rg, Nelson2025-zb, Stuhler2025-rp, Ashwin2025-af]. Furthermore, if sources of empirical data, such as open-ended survey responses [Zhang2025-st] or social media posts [sourati2025shrinking] are increasingly produced using GenAI, future research will be confronted with difficult questions about sociological explanation and inference.
@@ 3.1 Sample
We used a bibliometric multistage sampling design to construct a representative sample of both computational and non-computational sociologists and their collaborators. The sample includes 219 non-computational authors and an over-sample of 214 computational authors (defined as those authoring an article with computational terms in the title, abstract, and keywords) for a total of 433 respondents. We used rake weights [debell2009computing] to adjust for this oversampling as well as selection bias into our sample—namely, gender and location. Our sampling method is detailed in Appendix C. Figure 1 is a summary schematic of the sampling strategy.
@@ 3.2 Measures
Our survey focuses on who uses GenAI, its use cases, reasons for and concerns over using GenAI, and sources of optimism and trust in the technology. See the replication repository for information about accessing the survey questionnaire and data, which includes variables not explored in the main text.
@@ 3.3 Analytic Plan
One particular interest of our work is to describe differences between scholars who self-report using computational methods and those who do not. To do this, we use an array of univariate, bivariate, and multivariate descriptive data visualizations. We also use regression models to analyze the extent trust and optimism are associated with levels of GenAI use and familiarity. All descriptive analyses and regressions are rake-weighted with weight-adjusted standard errors and confidence intervals [lumley2017fitting]. See Appendix D for our findings as percentages over individual respondent categories (e.g., $X\%$ of respondents stated they “somewhat agree” with $Y$). Recall that our sample includes sociologists and their collaborators. To simplify, we refer to this population collectively as scholars.
@@ 4.2 GenAI Use Cases
Of those who have used GenAI in their research at least once, how do they use it? The second panel of Figure 2 shows the distribution of GenAI use at various stages of research—planning, writing, data collection, and analysis—separated by non-computational and computational scholars. The percentages sum to 100% per research stage, but with the never-users included in the calculation and removed from the visualization. As the bar chart shows, computational scholars are slightly more likely to use GenAI in analysis tasks (21.9%; 16% $\leq\mu\leq$ 29.2%) than non-computational scholars (13.6%; 8.4% $\leq\mu\leq$ 21.4%). Outside this research stage, however, the differences are marginal. Importantly, while the never-users are not shown, they make up the largest group for each research stage—in other words, most scholars do not use GenAI in any of the research stages as defined by the survey. Nonetheless, among those who have used GenAI, writing is the most common research stage where GenAI comes into play for both non-computational and computational scholars. In the Appendix (Figure 5 in Appendix D) we break down each research stage into specific tasks, such as translating one’s own writing or explaining statistical outputs. [^2] Generally, we find very few differences, the most significant of which is in the analysis stage. Specifically and unsurprisingly, computational scholars are using GenAI to help with coding – writing (20.8%) and debugging (21.8%). Non-computational scholars are also somewhat more likely to use GenAI for grammar and spelling assistance (32.6%) and translation (24.8%). That said, both computational and non-computational scholars most often use GenAI to help with writing tasks. Though analysis is tied with writing for computational scholars.
@@ 4.4 Concerns about GenAI
Overall, our items indicate that there is a high level of concern about various aspects of GenAI; of the nine questions we asked, seven were rated as being between “Concerned” and “Very Concerned” on average. To capture possible nuance regarding respondents’ perceptions of risks, we also asked about their GenAI concerns in an open-ended question. For instance, one respondent wrote that they were “wholly against the use of GenAI tools. They have no place in social science practice.” Below, we complement the aggregate findings with further examples that we found to be representative of general themes in the open-ended responses.
@@ 4.5 Optimism and Trust
Finally, we measure the optimism and trust about the future of GenAI for research. Scholars are generally more optimistic that GenAI will continue to improve over the next 2 to 3 years (40.2% of all scholars “Somewhat agree” or “Strongly agree”)—though there is ambivalence about whether this will have a net positive effect on the field or whether the current advantages of GenAI outweigh the drawbacks for the field. Furthermore, scholars generally believed that GenAI outputs cannot be trusted (only 4.5% of all scholars “Somewhat agree” or “Strongly agree” that GenAI can be trusted, see also Figure 10 in Appendix D).
@@
I treat GenAI as an inexperienced, mediocre RA. I don’t trust it, but use it to do simple things I can easily check
and
I think we’re at the early stages here. I think it can sometimes be useful to bounce ideas off, but often it returns quite derivative feedback, and I have not found these tools very useful for searching for articles.
We used rake-weighted generalized linear regression models to further probe heterogeneity in trust and optimism. [^3] We also ran a series of rake-weighted ordered logistic regression models as a robustness check given the ordinal nature of the outcome variables. They are consistent with the simpler linear model results reported here. These alternative models are in Table 6 in Appendix F. Adjusted predictions for these models are visualized in Figure 11. Specifically, we examine whether scholars who are more familiar with GenAI are more likely to trust GenAI, believe that GenAI will improve in the next 2 to 3 years, or believe that GenAI will have a net positive impact on their field in the next 2 to 3 years, and if those effects might be moderated by their use level.
@@
The regression results are in Table 5 in Appendix E with unstandardized estimates. A model with only familiarity as a predictor of trust (not shown in Table 5) had a very small positive effect ($\hat{\beta}$ = 0.078 with $t$ = 3.660). However, use level is not a significant moderator of this effect (see Model 1 in Table 5). Use level is also not a significant moderator for the net positive outcome variable but is for the future improvement variable. The regression results are summarized with adjusted predictions in Figure 4, where gender identity is held at its modal value (cis-man).
@@ 5 Discussion
Second, we do not yet know to what extent sociologists might see GenAI as a threat to their sociological expertise. Sociological work may be challenged by GenAI as even the specialties most removed from computational methodology, namely sociological theory, are implicated. [^5] Incidentally, the most elite sociology departments in the US specialize in theory [elder2025status]. More generally, GenAI threatens knowledge work by broadening the availability of required expertise, weakening occupational closure mechanisms, and reducing bargaining power. Such tensions are further exacerbated by the increased dependence on science and technology in tandem with a decreasing trust in scientists [eyal2019crisis]. This routinization of knowledge production may eventually allow for the displacement of well-paid professionals for “workers paid a fraction of their wage, flood[ing] the market with inferior products, and impos[ing] unreasonable and punishing working conditions on those” using GenAI [Bender2025-wa], c.f., [Nelson2025-zb]. Social scientists might come to feel these tensions more acutely than they do now, given that GenAI can present convincing or even compelling arguments about human society and social behavior, even if these are inaccurate or entirely fabricated.
@@
Appendix A Appendix: Survey Invitation Primer on GenAI
Appendix B Example Recruitment Materials
B.1 Invitation Email
@@ B.2 Generative AI Explainer
Appendix C Sampling Methods
C.1 Design
A scholar’s expertise, contributory or interactional à la Collins and Evans [collins2019rethinking], in relation to GenAI is likely to shape both usage and attitudes. Specifically, we expected that computational scholars would bring a potentially unique perspective on GenAI as it has emerged out of the same computational techniques that many such scholars use (i.e., machine learning and NLP). However, comparatively, this group is a very small subset of the broader sociological community. We therefore make an effort to both identify these scholars from our larger sampling frame and oversample them. In addition to having a better understanding of how these technologies operate, we also expected that computational scholars will have a higher response rate due to intrinsic interest and will be more likely to use GenAI in some capacity. Furthermore, based on prior surveys of (generative and non-generative) AI usage in academia, we expected that men will also be more likely to respond (possibly, in part, because they are overrepresented among computational scholars). We therefore design the survey to allow us to adjust for some parameters of our population. Although we attempt to control for such bias, inferences to the entire field of sociology must be made with caution. See the schematic representation of our sampling strategy in the main text(Figure 1).